# Tight Bounds for Sparsifying Random CSPs

### v1

@@ 7.6.1 Theorem Statement
Let $\mathbf{R}:D^{r}\to\mathbb{Z}_{\geq 0}$ be a valued relation which is precisely $k$-plentiful and marginally uniform. Then, for any $\varepsilon\in(0,1)$ the i.i.d. sparsifier of $C$ which preserves $\widetilde{O}_{\mathbf{R}}(n^{r-k}/\varepsilon^{3})$ constraints is a $(1\pm\varepsilon)$ sparsifier with high probability.
@@ 7.6.2 Sandwiching
First observe that any $h\in S_{\psi}$ with $h_{d}=k$ is $(d,k)$-tight, for if there is any $h^{\prime}\prec_{\bar{d}}h$ with $h^{\prime}_{d}\geq k+1$, then $h^{\prime}\in S_{\psi}$ showing that $\psi$ is not tight. By 7.18 $h$ is $(d,E)$-rigid. Therefore, since $\mathbf{R}$ is marginally uniform, for all $s\in D^{r}$ with $\operatorname{set}(s)=h$, for all $t\sim_{E}s$, we have that $\mathbf{R}(s)=\mathbf{R}(t)$. Thus, $\mathbf{R}_{E,0}(s)=\mathbf{R}_{E,1}(s)=\mathbf{R}(s)$, as desired. ∎
@@
Since $\operatorname{hist}(t^{\prime})$ is $(d,k)$-tight and $(d,E)$-rigid by 7.18, we can deduce that (1) $\operatorname{hist}(t^{\prime})\prec_{\bar{d}}\operatorname{hist}(s)$ as $\operatorname{hist}(t^{\prime})(E\setminus\{d\})=0$ and (2) since $\mathbf{R}$ is marginally uniform for all $s^{\prime}\sim_{E}t^{\prime}$ we have that $\mathbf{R}(s^{\prime})=\mathbf{R}(t^{\prime})$. Thus, we cannot have that $s\sim_{E}t^{\prime}$. In other words, there exists $e\in D\setminus E$ for which $_{e}t^{\prime}<\#_{e}s=\#_{e}t$.
@@ 7.7 Average case lower bound
The purpose of this section is to close one last gap in the analysis of random CSP instances. Assume that $\mathbf{R}$ is precisely $k$-plentiful and $\operatorname{Sym}(\mathbf{R})$ is marginally uniform, but $\mathbf{R}$ is not marginally uniform. What is the sparsifiability of a random instance with $\lambda\cdot n^{r-k}$ clauses, where $\lambda=O_{\mathbf{R}}(n)$? We show in this case no nontrivial sparsification can be achieved.
@@
Let $\mathbf{R}$ be precisely $k$-plentiful and not marginally uniform. Consider a random instance $I$ of $\operatorname{VCSP}(\mathbf{R})$ with $\lambda\cdot n^{r-k}$ clauses, where $\lambda=O_{\mathbf{R}}(n)$. Then any $(1\pm\varepsilon)$ sparsifier $\hat{I}$ of $I$ for $\varepsilon\in(0,O_{\mathbf{R}}(1))$ requires at least $\Omega_{\mathbf{R}}(\lambda\cdot n^{r-k})$ edges–even if $\hat{I}$ is not a subset of $I$.
@@
Since $\mathbf{R}$ is not marginally uniform, there is $d\in E\subseteq D$ such that $h\in\mathsf{T}_{d,k}(\operatorname{Sym}(\mathbf{R}))$ is $(d,E)$-rigid as well as $s,t\in D^{r}$ with $s\sim_{E}t$, $\operatorname{set}(s)=h$, and $\mathbf{R}(s)\neq\mathbf{R}(t)$.
@@ 8 Translating Sparsifiability Bounds to Random Instances
    If $R$ is precisely $k$-plentiful, and the symmetrization of $R$ is not marginally uniform, then (1) $C$ is sparsifiable to $\widetilde{O}_{\mathbf{R}}(\min(m,n^{r-k+1}/\varepsilon^{2}))$ constraints, and (2) for a sufficiently small, constant $\varepsilon$, any sparsifier of $C$ requires $\Omega_{\mathbf{R}}(\min(m,n^{r-k+1}))$ constraints.
@@
    If $R$ is precisely $k$-plentiful, and $R$ is marginally uniform, then (1) $C$ is sparsifiable to $\widetilde{O}_{\mathbf{R}}(\min(m,n^{r-k}/\varepsilon^{3}))$ constraints, and (2) for a sufficiently small, constant $\varepsilon$, any sparsifier of $C$ requires $\Omega_{\mathbf{R}}(\min(m,n^{r-k}))$ constraints.
@@
    If $R$ is precisely $k$-plentiful, the symmetrization of $R$ is marginally uniform, but $R$ itself is not marginally uniform, then (1) if $m\geq\kappa(\mathbf{R})\cdot n^{r-k+1}/\varepsilon^{2}$ for a sufficiently large constant $\kappa(\mathbf{R})$, $C$ is $(1\pm\varepsilon)$-sparsifiable to $\widetilde{O}_{\mathbf{R}}(n^{r-k}/\varepsilon^{3})$ constraints, but (2) if $m\leq n^{r-k+1}/\kappa^{\prime}(\mathbf{R})$ for a sufficiently large $\kappa^{\prime}(\mathbf{R})$, there is a small, constant $\varepsilon$, such that any sparsifier of $C$ requires $\Omega_{\mathbf{R}}(m)$ constraints.
@@ 8.1 Bounds for Random Uniform Instances
Next, if $m=O_{\mathbf{R}}(n^{r-k+1})$, then for a sufficiently large choice of $\kappa^{\prime}(\mathbf{R})$, Theorem 7.28 immediately tells us that there exists a constant $\varepsilon$ such that any $(1\pm\varepsilon)$ sparsifier of $C$ must preserve $\Omega(m)$ re-weighted constraints. Otherwise, if $m=\Omega_{\mathbf{R}}(n^{r-k+1})$, by the above, there is a choice of $\ell=O_{\mathbf{R}}(n^{r-k+1}/\varepsilon^{2})$, such that a random sample of $m\geq\ell$ constraints (with weight $\binom{n}{r}\cdot r!/m$ each), is a $(1\pm\varepsilon/10)$ sparsifier of the complete CSP with high probability (at least $9/10$). In particular, any $(1\pm\varepsilon/10)$ sparsifier of $C$ would be (when weighted by $\binom{n}{r}\cdot r!/m$), a $(1\pm\varepsilon/10)^{2}$ sparsifier, and thus a $(1\pm\varepsilon)$ sparsifier of the complete CSP. But, by Theorem 7.15, Item 4, there exists a choice of $\varepsilon=\Omega_{\mathbf{R}}(1)$ such that any $(1\pm\varepsilon)$ sparsifier of the complete instance requires $\Omega_{\mathbf{R}}(n^{r-k+1})$ constraints. ∎
@@
If $m\geq\kappa(\mathbf{R})\cdot n^{r-k+1}/\varepsilon^{2}$ for a sufficiently large constant $\kappa(\mathbf{R})$, then, as in Lemma 7.6, we can observe that $C$ re-weighted by $\frac{\binom{n}{r}\cdot r!}{m}$ is a $(1\pm\varepsilon/10)$ sparsifier of the complete, uniform CSP instance, as these sparsifiers are created by uniform, random sampling with replacement. In particular, we can then invoke Theorem 7.15, Item 1 which states that if $\operatorname{Sym}(\mathbf{R})$ is marginally uniform, then for any $\varepsilon\in(0,1)$, there exists a $(1\pm\varepsilon)$ sparsifier (denote this $\hat{C}$) of the complete instance which preserves $\widetilde{O}_{\mathbf{R}}(n^{r-k}/\varepsilon^{3})$ constraints. Now, by composing sparsifiers, we obtain that $\hat{C}$ is a $(1\pm\varepsilon/10)^{2}$ and thus a $(1\pm\varepsilon)$ sparsifier of our random CSP $C$ (re-weighted by $\frac{\binom{n}{r}\cdot r!}{m}$). Multiplying by the inverse of the weighting then yields the upper bound.

### v2

@@ 7.6.1 Theorem Statement
Let $\mathbf{R}:D^{r}\to\mathbb{Z}_{\geq 0}$ be a valued relation which is precisely $k$-plentiful and marginally balanced. Then, for any $\varepsilon\in(0,1)$ the i.i.d. sparsifier of $C$ which preserves $\widetilde{O}_{\mathbf{R}}(n^{r-k}/\varepsilon^{3})$ constraints is a $(1\pm\varepsilon)$ sparsifier with high probability.
@@ 7.6.2 Sandwiching
First observe that any $h\in S_{\psi}$ with $h_{d}=k$ is $(d,k)$-tight, for if there is any $h^{\prime}\prec_{\bar{d}}h$ with $h^{\prime}_{d}\geq k+1$, then $h^{\prime}\in S_{\psi}$ showing that $\psi$ is not tight. By 7.18 $h$ is $(d,E)$-rigid. Therefore, since $\mathbf{R}$ is marginally balanced, for all $s\in D^{r}$ with $\operatorname{set}(s)=h$, for all $t\sim_{E}s$, we have that $\mathbf{R}(s)=\mathbf{R}(t)$. Thus, $\mathbf{R}_{E,0}(s)=\mathbf{R}_{E,1}(s)=\mathbf{R}(s)$, as desired. ∎
@@
Since $\operatorname{hist}(t^{\prime})$ is $(d,k)$-tight and $(d,E)$-rigid by 7.18, we can deduce that (1) $\operatorname{hist}(t^{\prime})\prec_{\bar{d}}\operatorname{hist}(s)$ as $\operatorname{hist}(t^{\prime})(E\setminus\{d\})=0$ and (2) since $\mathbf{R}$ is marginally balanced for all $s^{\prime}\sim_{E}t^{\prime}$ we have that $\mathbf{R}(s^{\prime})=\mathbf{R}(t^{\prime})$. Thus, we cannot have that $s\sim_{E}t^{\prime}$. In other words, there exists $e\in D\setminus E$ for which $_{e}t^{\prime}<\#_{e}s=\#_{e}t$.
@@ 7.7 Average case lower bound
The purpose of this section is to close one last gap in the analysis of random CSP instances. Assume that $\mathbf{R}$ is precisely $k$-plentiful and $\operatorname{Sym}(\mathbf{R})$ is marginally balanced, but $\mathbf{R}$ is not marginally balanced. What is the sparsifiability of a random instance with $\lambda\cdot n^{r-k}$ clauses, where $\lambda=O_{\mathbf{R}}(n)$? We show in this case no nontrivial sparsification can be achieved.
@@
Let $\mathbf{R}$ be precisely $k$-plentiful and not marginally balanced. Consider a random instance $I$ of $\operatorname{VCSP}(\mathbf{R})$ with $\lambda\cdot n^{r-k}$ clauses, where $\lambda=O_{\mathbf{R}}(n)$. Then any $(1\pm\varepsilon)$ sparsifier $\hat{I}$ of $I$ for $\varepsilon\in(0,O_{\mathbf{R}}(1))$ requires at least $\Omega_{\mathbf{R}}(\lambda\cdot n^{r-k})$ edges–even if $\hat{I}$ is not a subset of $I$.
@@
Since $\mathbf{R}$ is not marginally balanced, there is $d\in E\subseteq D$ such that $h\in\mathsf{T}_{d,k}(\operatorname{Sym}(\mathbf{R}))$ is $(d,E)$-rigid as well as $s,t\in D^{r}$ with $s\sim_{E}t$, $\operatorname{set}(s)=h$, and $\mathbf{R}(s)\neq\mathbf{R}(t)$.
@@ 8 Translating Sparsifiability Bounds to Random Instances
    If $R$ is precisely $k$-plentiful, and the symmetrization of $R$ is not marginally balanced, then (1) $C$ is sparsifiable to $\widetilde{O}_{\mathbf{R}}(\min(m,n^{r-k+1}/\varepsilon^{2}))$ constraints, and (2) for a sufficiently small, constant $\varepsilon$, any sparsifier of $C$ requires $\Omega_{\mathbf{R}}(\min(m,n^{r-k+1}))$ constraints.
@@
    If $R$ is precisely $k$-plentiful, and $R$ is marginally balanced, then (1) $C$ is sparsifiable to $\widetilde{O}_{\mathbf{R}}(\min(m,n^{r-k}/\varepsilon^{3}))$ constraints, and (2) for a sufficiently small, constant $\varepsilon$, any sparsifier of $C$ requires $\Omega_{\mathbf{R}}(\min(m,n^{r-k}))$ constraints.
@@
    If $R$ is precisely $k$-plentiful, the symmetrization of $R$ is marginally balanced, but $R$ itself is not marginally balanced, then (1) if $m\geq\kappa(\mathbf{R})\cdot n^{r-k+1}/\varepsilon^{2}$ for a sufficiently large constant $\kappa(\mathbf{R})$, $C$ is $(1\pm\varepsilon)$-sparsifiable to $\widetilde{O}_{\mathbf{R}}(n^{r-k}/\varepsilon^{3})$ constraints, but (2) if $m\leq n^{r-k+1}/\kappa^{\prime}(\mathbf{R})$ for a sufficiently large $\kappa^{\prime}(\mathbf{R})$, there is a small, constant $\varepsilon$, such that any sparsifier of $C$ requires $\Omega_{\mathbf{R}}(m)$ constraints.
@@ 8.1 Bounds for Random Uniform Instances
Next, if $m=O_{\mathbf{R}}(n^{r-k+1})$, then for a sufficiently large choice of $\kappa^{\prime}(\mathbf{R})$, Theorem 7.28 immediately tells us that there exists a constant $\varepsilon$ such that any $(1\pm\varepsilon)$ sparsifier of $C$ must preserve $\Omega(m)$ re-weighted constraints (note that we can invoke Theorem 7.28 because if $\mathrm{Sym}(\mathbf{R})$ is not marginally balanced, then $\mathbf{R}$ is not marginally balanced). Otherwise, if $m=\Omega_{\mathbf{R}}(n^{r-k+1})$, by the above, there is a choice of $\ell=O_{\mathbf{R}}(n^{r-k+1}/\varepsilon^{2})$, such that a random sample of $m\geq\ell$ constraints (with weight $\binom{n}{r}\cdot r!/m$ each), is a $(1\pm\varepsilon/10)$ sparsifier of the complete CSP with high probability (at least $9/10$). In particular, any $(1\pm\varepsilon/10)$ sparsifier of $C$ would be (when weighted by $\binom{n}{r}\cdot r!/m$), a $(1\pm\varepsilon/10)^{2}$ sparsifier, and thus a $(1\pm\varepsilon)$ sparsifier of the complete CSP. But, by Theorem 7.15, Item 4, there exists a choice of $\varepsilon=\Omega_{\mathbf{R}}(1)$ such that any $(1\pm\varepsilon)$ sparsifier of the complete instance requires $\Omega_{\mathbf{R}}(n^{r-k+1})$ constraints. ∎
@@
If $m\geq\kappa(\mathbf{R})\cdot n^{r-k+1}/\varepsilon^{2}$ for a sufficiently large constant $\kappa(\mathbf{R})$, then, as in Lemma 7.6, we can observe that $C$ re-weighted by $\frac{\binom{n}{r}\cdot r!}{m}$ is a $(1\pm\varepsilon/10)$ sparsifier of the complete, uniform CSP instance, as these sparsifiers are created by uniform, random sampling with replacement. In particular, we can then invoke Theorem 7.15, Item 1 which states that if $\operatorname{Sym}(\mathbf{R})$ is marginally balanced, then for any $\varepsilon\in(0,1)$, there exists a $(1\pm\varepsilon)$ sparsifier (denote this $\hat{C}$) of the complete instance which preserves $\widetilde{O}_{\mathbf{R}}(n^{r-k}/\varepsilon^{3})$ constraints. Now, by composing sparsifiers, we obtain that $\hat{C}$ is a $(1\pm\varepsilon/10)^{2}$ and thus a $(1\pm\varepsilon)$ sparsifier of our random CSP $C$ (re-weighted by $\frac{\binom{n}{r}\cdot r!}{m}$). Multiplying by the inverse of the weighting then yields the upper bound.
@@ 8.2 Bounds for Random $r$-Partite Instances
Note that using the same reasoning above, but instead with Theorem 6.4 yields an analogous characterization for $r$-partite valued CSPs, yielding Theorem 1.4.